Cycle's wheel radius is 45 cm. Find the number of it's turnings while riding from school to home if distance between is 1 km.
1
Expert's answer
2010-06-16T06:16:43-0400
During one turning wheel passes the following distance: l=2*pi*R=2*3.14*0.45=2.826 m,
so if the distance from home to school is 1 km we can find the number of turnings by dividing distance between school and home by the distance, which wheel passes during one turn: 1000/2.826=353,8